OpenStudy (anonymous):

What is the mass of 5.25 mols of boron trifluoride?

OpenStudy (anonymous):

356 grams

OpenStudy (anonymous):

Could you explain to me how you did that, please?

OpenStudy (anonymous):

mass/molecular weight of BF3 = moles mass/67.805=5.25 mass=356grams mw of BF3= mw of B + mw of 3(F) = 10.811 + 3(18.988) =10.811+56.94 = 67.85